OpenStudy (anonymous):

Find the domain of the function. f(x) = -2x + 4

OpenStudy (anonymous):

A. (-∞, 0) (0, ∞) B. (-∞, ∞) c.[-4, ∞) D.(0, ∞)

OpenStudy (triciaal):

the domain in the set of x values

OpenStudy (anonymous):

I done one of these while ago, and got it wrong

OpenStudy (anonymous):

okay..well,my teacher called it mother function :D

OpenStudy (triciaal):

what is f(0)?

OpenStudy (anonymous):

0?

OpenStudy (triciaal):

when you let x = 0 what is f(x)?

OpenStudy (triciaal):

f(x) = -2x + 4 f(0) = -2(0) + 4 = 0 + 4 =

OpenStudy (anonymous):

i gree with you @triciaal

OpenStudy (anonymous):

4?

OpenStudy (triciaal):

4? or 4.

OpenStudy (anonymous):

2 and 4

OpenStudy (triciaal):

0 + 4 = 4 I want you to make a statement. be sure on the little things

OpenStudy (triciaal):

now what is x when f(x) = 0 this means solve for x when -2x + 4 = 0

OpenStudy (anonymous):

-2x+4 = 4

OpenStudy (triciaal):

no read again

OpenStudy (triciaal):

@ciinpi "okay..well,my teacher called it mother function :D" what does this mean? @kaykayrae89 -2x + 4 = 0 solve for x

OpenStudy (anonymous):

-2(2)=4=0 x=2

OpenStudy (triciaal):